An 8-year annuity due has a present value of $1,000. If the interest rate is 5 percent, what is the amount of each annuity payment?
Answer #1

Ans $ 154.72

Annuity PV Factor (End of Year) = P [ 1 - ( 1 + r )^-n ] / r

Annuity PV Factor (End of Year) = P[1-(1+r)^-n] 1000 = P*(1-((1/(1+5%)^8) 5% 1000 = 0.323160637971313* P 0.05 P= 1000 * 0.05